1) phosphorus pentachloride is produced by the reaction between phosphine and chlorine gas. What mass of chlorine gas would be required to produce 0.015kg of phosphorus pentachloride
2) if 10g of phosphine is made to react with 15g of chlorine gas. Calculate the percentage yield if 5.5g of phosphorus pentachloride was experimentally produced
1
Expert's answer
2019-04-01T07:48:30-0400

2 PH3 + 8 CL2 = 2 PCL5 + 6 HCL

M(Cl2) = 71 g/mol

M(PCL5) = 208.24 g/mol

M(PH3) = 40.0 g/mol

1) For equation :

568 g (CL2) = 416.48 g (PCL5)

x g (Cl2) = 15 g (PCL5)

x = 568*15/416.48 = 20.46 g (CL2)

Answer : 20.46 g CL2

2) For equation we can see, that PH3 is an excess. So the calculations are carried out on chlorine.

For equation :

568 g (CL2) = 416.48 g (PCL5)

15 g (Cl2) = x g (PCL5)

x = 416.48*15/568 = 11.0g (PCL5)

For experiment yield is 5.5g

Percentage yield = 5.5/11.0*100% = 50%

Answer : 50%